13.) Assume y varies inversely as x. If y = -32 when x = 2, what is x when y = 4?
1
Expert's answer
2012-05-11T07:47:02-0400
y-new/y_old=4/(-32)=-1/8
x_new/x_old=-8 => x_new=-8*2=-16